Present: Brian Burton, Bill Covey, Debra Harris, Winston Harris, Ashley Jones, Everett Phillips, Ying TangAnnouncements
A. Security Problems - theft of 2 monitors from the Systems vault will result in new procedures:
B. Signage - Rob Roberts reported at the last MMM meeting that he has received quite a few complaints about the information & placement of signs around the libraries. The signs were recently installed by a contracted, commercial vendor. We have requested the removal of the Systems name from the Preservation door & that the label at our front door be less cryptic.
- Supplies - supplies will be moved to the closet located across from Bill's office. There will no longer be any reason to leave the vaults open.
- Vaults - we will try to get the combination on both vaults changed. The numerical combinations are not to be written down on any piece of paper and kept in the office. The only people who will need to know the combination will be Bill and the technical staff.
- Visitor Log - anyone setting foot in Systems will need to sign-in at our front desk. This includes all library personnel, PPD, housekeeping, recycle workers, Culligan water delivery personnel, fire extinguisher inspectors and even the person delivering the mail (if they do not use an intellikey to enter the department).
- Other - we have the option of installing a buzzer on the doors so that they will buzz when people leave the building. We could install 24-hour video cameras in the vaults. Could set the cameras to come on when detecting motion or when the doors open. Would have to wait & see if the money would be made available. We do not want to get the reputation of being an easy mark for thieves.
C. Whistling valve problem - Bill will try to have the screech reduced.
Review Minutes of 6/7/99
- Announcement #3 - the rumor that campus is off ATM is supposedly true.
Ongoing Systems Updates
- Three Dell servers received. Winston upgraded SmathersNT4 with one of them. The process took 14 hours over the weekend, hopefully this will decline with the next two installations. The server is a super machine - a step up from the mail server. The biggest problem Winston encountered was locating the raid & SCSI drivers. Tech support was called for assistance and eventually the server was up & running. Winston got the print queues up this morning. Winston switched the SQL server to NT security. Staff will not need to enter a password if their SQL access is on NT4. The power supplies are loud - may eventually move them into Grace's office. NT3 may move there as well, which will provide for more security. Plans for the other two servers: one is to go in the North tower to service Resource Services, LAC & Preservation on the 223 subnet. This way we'd still have access to the files in the print queues if the rest of campus went down. The final server is currently destined to upgrade NT5 in MSL, but given the uncertainty of the Preservation split - it may have to go in the North tower as well - then NT5 will be upgraded at a later date.
- No word on the Cow shipment.
- 45 switches are due this Friday. The goal is have everyone on switched ports, replacing the hubs.
- Not much more will be ordered until after July 1st.
- Martha reported that the mail Bill sent her was immediately disappearing out of her mailbox. Bill will check on it. Spams & viruses abound.
- Current Wiring - Bill & Ashley will be checking on the progress with a walk-through today.
- Bookstore on the second floor is not ready yet, but Digital is scheduled to do the punchdowns this week. This is not a critical project.
- Documentation project to move into high gear. Will need it completed by Friday to accommodate the incoming switches.
Software Projects Status
- Personnel database - Winston will meet with Carolyn tomorrow.
- SQL server - Ying's sign-on via NT4 is giving her an inconsistent level of service: sometimes she can edit, sometimes not. Microsoft has been contacted, reinstallation was performed - but the problem still exists.
- LASSIE - had a glitch, but it's working now.
- Mellon - new release is out with an updated Cuban Missile Crisis section.
- Compustat tapes were mailed to the University of Iowa. We are awaiting word from them.
- The break between summer A & B sessions is next week, June 21-25. Some students have expressed interest in working.
- Suzy will be out all week at the InterDev Conference.
- Rich out today.
- Ashley's last day is Tuesday, June 15.
- Everett out June 21.
- Rich out June 25-29 (ALA).
- Winston out July 1 & 2.
- Everett will be out July 2-9.
- Bobbie, Ying and Winston out July 12-16.
- Brian will not be out July 30.
Back to the Systems Homepage
Last updated February 18, 2000
by Debra Harris